OpenStudy (anonymous):

What is the common difference in this sequence? 10, 17, 24, 31,... A. 10 B. 7 C. 21 D. 3.1

Parth (parthkohli):

This is an arithmetic progression and it has a common difference of 7. It keeps adding 7 each time.

OpenStudy (maheshmeghwal9):

b

Parth (parthkohli):

@maheshmeghwal9 You should really explain your ways. Giving the answer can suspend you. Okay?

OpenStudy (maheshmeghwal9):

sorry!

OpenStudy (maheshmeghwal9):

10+7=17=>17+7=24=>24+7=31...... & so on.!.:)